# Leveraging Graph Neural Networks and Multi-Agent Reinforcement Learning for Inventory Control in Supply Chains

### v1

@@ 3.3 Problem Statement
Moreover, the action space is traditionally modelled as the order replenishment quantity, $o$. However, in this paper we parameterize a heuristic inventory policy such that for each agent $i$, at each time step $t$, the policy outputs a mean and standard deviation for both $s_{\text{inv},t}^{i}$ and $S_{\text{ord},t}^{i}$. Therefore, $\pi_{i}(o^{i})=(\mu_{s_{\text{inv},t}^{i}},\sigma_{s_{\text{inv},t}^{i}},\mu_{S_{\text{ord},t}^{i}},\sigma_{S_{\text{order},t}^{i}})$ and the heuristic inventory parameters are sampled from a Gaussian distribution, $s_{\text{inv},t}^{i}\sim\mathcal{N}(\mu_{s_{\text{inv},t}^{i}},\sigma_{s_{\text{inv},t}^{i}}^{2})$, $S_{\text{ord},t}^{i}\sim\mathcal{N}(\mu_{S_{\text{ord},t}^{i}},\sigma_{S_{\text{ord},t}^{i}}^{2})$. A min-max post processing step is then used to scale the values to a suitable range denoted by a subscript $s$, leading to $(s_{t_{\text{inv,s}}}^{i},S_{t_{\text{ord,s}}}^{i})$ where $s_{t_{\text{inv,s}}}^{i}$ is the reorder point and $S_{t_{\text{ord,s}}}^{i}$ is the order-up-to-level. When the inventory reaches a level of $s_{t_{\text{inv,s}}}^{i}$, an inventory order is placed where $o_{r,t}^{i}=S_{t_{\text{ord,s}}}^{i}-v_{t}^{i}$. The neural network architecture for each actor, including the post-processing step that results in the order replenishment quantity, is illustrated in Figure 5.
@@
Moreover, in inventory management, optimal order quantity policies such as $(s_{t_{\text{inv,s}}}^{i},S_{t_{\text{ord,s}}}^{i})$ are often characterized by discrete functions as they show abrupt changes in order quantity at specific inventory levels Dehaybe et al. (2024). This poses a challenge for neural networks, which are inherently continuous function approximators as directly approximating such discrete policies can lead to instability and poor performance in neural network-based models. To address this, in this paper, the actor network outputs a normal distribution (Gaussian) over the $(s_{t_{\text{inv}}}^{i},S_{t_{\text{ord}}}^{i})$ parameters. A key note is that this approach does not limit the network’s ability to learn the optimal policy as it simply expresses the policy in a way that may be easier to learn. Moreover, unlike traditional heuristic methods where the policy is not conditioned on the state of the system, an RL approach means the heuristic $(s_{t_{\text{inv,s}}}^{i},S_{t_{\text{ord,s}}}^{i})$ policy is conditioned on the state of the system such as on-hand, pipeline inventory and backlog. This enables the agent to dynamically adjust the reorder point and order-up-to level based on the current inventory situation, potentially leading to optimal and flexible decision-making.
Finally, while the actual order replenishment quantities in our environment are discrete, we model the action space for each agent as continuous within the range of $\left[-1,1\right]$. This approach facilitates scalability to problems with a wider range of possible order sizes. To bridge the gap between the continuous action space and the discrete nature of order quantities, we employ a post-processing step using min-max scaling. During this step, the agent’s chosen action is scaled to the appropriate range for the order replenishment variable and then rounded to the nearest integer value.
Figure 6: Schematic illustrating the execution of the trained policies simultaneously alongside heuristic inventory control policies.
@@ 4 Results and Discussion
In this section four case studies of different inventory management configurations are used to illustrate the effectiveness of our proposed methodology as shown in Figure 7. We compare our proposed methodology (Noisy P-GCN-MAPPO and P-GCN-MAPPO) against other MARL methods including: IPPO, MAPPO, and Graph-based-MAPPO(G-MAPPO). This allows us to compare against state-of-the-art as well as analyse the contribution of the different components of our method. One of the key differences between the methods is the varying levels of information input to the value function which can be illustrated qualitatively in Figure 8. This is further benchmarked against single agent RL (PPO specifically) and a heuristic (s,S) policy. In the later, a static heuristic policy is found for each node in the network, where optimal parameters are found using a derivative-free method, implemented using SciPy Virtanen et al. (2020) with a multi-start approach combined with local search.
All variant MARL algorithms are implemented in the Ray RLLib framework Liang et al. (2018), and all the hyperparameters are kept the same as the baseline algorithms to ensure fairness.
@@
Figure 8: Qualitative representation of varying levels of information input to the value function. The different levels illustrate how changes in observability can affect performance.
@@ 4.1 Execution
Firstly, the bar graph shown in Figure 9 compares the cumulative profit for the different methodologies in the different supply chain network configurations. Notably, the proposed methodology, Noisy P-GCN-MAPPO, has on par performance or outperforms the other methodologies. This advantage becomes increasingly pronounced as the number of agents rises. This trend suggests that Noisy P-GCN-MAPPO may be particularly effective in large-scale, multi-agent inventory control scenarios. The decrease in cumulative profit from the 18 agent to 24 agent configuration can be attributed to differences in number of retailers present and the specific environment configuration chosen. Figure 9 also shows that while the static $(s_{t_{\text{inv,s}}}^{i},S_{t_{\text{ord,s}}}^{i})$ policy achieves similar cumulative profit performance to Noisy P-GCN-MAPPO, it exhibits a significantly higher standard deviation. This variability suggests that the static heuristic is less stable, potentially hindering its robustness in situations with unexpected changes to the environment.
This distinction is crucial in real-world applications, where supply chain networks and other multi-agent systems inevitably encounter unexpected events. RL’s ability to maintain good performance in uncertain environments is one of the key advantages compared to traditional methods. Therefore, the results highlight the strength of RL methods, particularly Noisy P-GCN-MAPPO, as its lower standard deviation suggests greater potential for robustness and adaptability in uncertain environments.
@@
Our methodology is also compared with other MARL methods. The notable difference between them is the information available to the critic. One key challenge in MARL is balancing access to global information (mitigating non-stationarity) with scalability limitations of naively combining all agent states. Our proposed methodologies, P-G-MAPPO and Noisy P-GCN-MAPPO, address this by leveraging graph-based approaches and the inherent structural properties of the problem. This allows agents to exploit both local and global information effectively.
@@
Figure 11 shows how IPPO consistently outperforms MAPPO, and this performance gap widens with a greater number of agents. This motivated the development of smarter methods of information aggregation, by leveraging graph-based approaches, where information can be harnessed whilst not suffering from performance shortcomings. This highlights the need for smarter information aggregation. G-MAPPO, P-GCN-MAPPO and Noisy P-GCN-MAPPO incorporate Graph Neural Networks (GNNs) to enable communication and information sharing while maintaining focus on local details.
@@
However, from Figure 11, the G-MAPPO method does not always outperform MAPPO and IPPO methods. This may occur for similar reasons why IPPO outperforms MAPPO. Too much information may be captured leading to the policies overfitting. Therefore, P-GCN-MAPPO outperforms G-MAPPO, MAPPO and IPPO in all 4 case studies. This occurs as the integration of a pooling mechanism, can help aggregate information from different parts of the graph, reducing the dimensionality of the data acting as a regularizer and potentially reducing overfitting. The message passing inherent in graph neural networks allows for agents to inherently communicate with each other whilst the pooling mechanism prevents overfitting of the policies. Despite losing local information in the critic, the global mean pooling mechanism allows for global information to be captured in the central critic that helps reduce the variance in the individual actors that are trained with local information.
@@
Our final proposed methodology hypothesizes that the addition of noise into the value function, propagates variability into the advantage function potentially reducing overfitting. The execution curves shown in Figure 11 show the impact of adding noise to the value function on our methodology’s performance. The results show that with a small number of agents (6 agents), adding noise into the value function does not outperform P-GCN-MAPPO compared to when the number of agents is larger (24 agents). When the number of agents is 6, the overall state space is smaller compared to a large agent system. Therefore, the addition of noise will just add unnecessary randomness which could disrupt the learning process as opposed to benefit the training of the policies. However, as the number of agents increases, it becomes harder for traditional MARL algorithms to explore the entire space efficiently. Therefore, large agent systems are more prone to getting stuck in local optima. The addition of noise to the value function encourages the agents to explore a wider range of actions, leading towards better solutions. The addition of noise can also be seen as regularization, where the policies become more robust to the complex environment.
@@
Moreover, it is useful to compare the change in entropy during the training phase for each of the different methods. In the context of MARL, entropy quantifies the uncertainty or randomness in the policy’s action distribution (Sutton, 2018). During the training process, as the agents learn, the entropy in the policy distribution decreases as the agents learn better policies. Figure 12 shows that the rate of entropy decrease differs between the different methods. The addition of noise to the value function introduces controlled variability which propagates into the advantage estimates. This helps maintain a higher entropy for a longer period during training which is desirable as high entropy promotes exploration. During MARL training, agents need to balance exploring new actions and exploiting known good actions. By exploring a wider range of actions (due to high entropy), agents are more likely to discover better strategies and avoid getting stuck in sub-optimal solutions. This can be especially important in complex environments where the best course of action depends on what other agents are doing. Figure 12 shows that Noisy P-GCN-MAPPO, maintains a higher entropy throughout training, promoting exploration and reducing policy overfitting leading to superior performance.
Figure 12: Figure shows the change in entropy during the training process for the different methods in the 18 agent configuration

### v2

@@ 3.1 Problem Statement
Moreover, in inventory management, optimal order quantity policies such as $(s_{t_{\text{inv,s}}}^{i},S_{t_{\text{ord,s}}}^{i})$ are often characterized by discrete functions as they show abrupt changes in order quantity at specific inventory levels (Dehaybe et al., 2024). This poses a challenge for neural networks, which are inherently continuous function approximators as directly approximating such discrete policies can lead to instability and poor performance in neural network-based models. To address this, in this paper, the actor network outputs a normal distribution (Gaussian) over the $(s_{t_{\text{inv}}}^{i},S_{t_{\text{ord}}}^{i})$ parameters. A key note is that this approach does not limit the network’s ability to learn the optimal policy as it simply expresses the policy in a way that may be easier to learn. Moreover, unlike traditional heuristic methods where the policy is not conditioned on the state of the system, an RL approach means the heuristic $(s_{t_{\text{inv,s}}}^{i},S_{t_{\text{ord,s}}}^{i})$ policy is conditioned on the state of the system such as on-hand, pipeline inventory and backlog. This enables the agent to dynamically adjust the reorder point and order-up-to level based on the current inventory situation, potentially leading to optimal and flexible decision-making.
3.2 Graph Representation of Supply Chain Systems
Graph-based systems are commonly used in various domains to model complex networks of interactions. For example, transportation networks, social networks, recommendation systems, and communication networks all rely on graph representations to model the connections and relationships between different entities. Similarly, in supply chain management, viewing the system as a graph allows us to better understand and optimize the flow of goods, information, and resources across interconnected entities.
A supply chain can naturally be represented as a graph $G=(W,E)$ where the entities within the supply chain are represented as nodes $i\in W$, and the relationships or interactions between these entities are represented as edges $E$. This graph representation provides a framework for modeling the complex interactions and dependencies that characterize inventory management systems.
To effectively apply this graph-based approach, it is essential to translate a real-world supply chain into the corresponding graph representation. The key components of this translation are outlined below:
Nodes (Entities) Each node $i\in W$ corresponds to an agent responsible for managing inventory at that entity’s location where $W$ represents different entities in the supply chain..
Nodes Features Each node $i$ has a node feature $x^{i}\in X$ which represents the relevant observations for the agent located at that node. These features capture the critical variables needed for decision-making at each entity. In the context of inventory management, these features include the current inventory level, backlog, pipeline inventory, historical demand and order history.
Edges The edges $E$ in the graph represent the relationships of interactions between the different entities. These relationships include direct transportation links, supply routes, or communication channels between entities. In this work, an edge exists between two nodes if they are connected and if the flow of goods or orders is permitted between them.
Neighborhood The neighborhood of a node $i$, denoted by $M(i)$ is defined as the set of neighboring nodes $j$ connected to node $i$ via edges in $E$. Mathematically, this can be represented as $M(i)=\{j|(i,j)\in E\}$. These neighboring nodes represent other entities that directly interact with the current node in the supply chain. For each node $i$, the features of its neighbors $j$ are represented as $x_{j}^{i}$, capturing the relevant observations of the neighboring agents and providing the context for decision-making at node $i$.
This graph-based representation provides a flexible framework that can be applied to a wide variety of supply chain settings. By modeling the supply chain as a graph, we can capture the complex interactions between entities, allowing us to further improve collaboration when coordinating decisions across different agents.
3.3 Graph Convolutional Networks (GCNs) combined with Multi-Agent Proximal Policy Optimization (MAPPO) and a Pooling strategy (P-GCN-MAPPO)
In environments where the graph structure can be leveraged, Graph Neural Networks (GNNs) are commonly integrated into RL frameworks. GNNs were developed to efficiently leverage the structure and properties of graphs. GNNs operate on graph-structured data and are able to capture complex relationships and dependencies inherent in graphs. In this work, we use Graph Convolutional Networks (GCNs) combined with Multi-Agent Proximal Policy Optimization (MAPPO) and a Pooling strategy, hence P-GCN-MAPPO. In Section 4 we conduct computational experiments to analyze the different components of this methodology in an effort to distill their contribution to the overarching framework. GCNs update the representation of a node by aggregating and transforming the features of its neighbouring nodes and itself. This allows the model to capture and propagate local information, effectively learning patterns and dependencies from the graph structure of a supply chain.
Traditional model-free methods rely solely on agent-environment interactions, whereas our GNN-based framework incorporates geometric information from the supply chain’s topology to enhance learning (Almasan et al., 2022; Zhou et al., 2020a). While standard model-free approaches can struggle with capturing complex dependencies, Graph Neural Networks (GNNs) leverage structural relationships to learn richer representations, even in the absence of a perfect system model (Yang et al., 2022). This ability makes GNNs well-suited for supply chain problems, which are characterized by inherent uncertainties and complex interactions. Moreover, GNNs are robust to noisy and imperfect data, effectively capturing the underlying relationships in the presence of missing or uncertain information (Verma et al., 2021; Jin et al., 2021). This ability to capture complex relationships is particularly beneficial for supply chain problems as understanding the intricate relationships between various entities can significantly optimize decision-making.
The adjacency matrix, $A\in\mathbb{R}^{N\times N}$, is a $N\times N$ matrix which is used to express the directed graph topology where $N$ is the number of nodes (or vertices) in the graph. In this matrix, $A_{ij}=1$ indicates that there is an edge between node $i$ and node $j$. Each node $i$ is associated with a node feature $x_{i}$ as described in Section 3.2, which encapsulates information specific to that node where $x_{i}\in\mathbb{R}^{D}$ where $D$ represents the dimensionality of the features vector. These individual node features collectively form the node feature matrix, $X\in\mathbb{R}^{N\times D}$. At each time step, the node feature matrix, $X$, captures the evolving state of the graph. Both $A$ and $X$ are fed into a graph convolution layer, allowing the model to capture relational information between nodes. The function $f(X,A)$ represents the graph convolution operator which aggregates and transforms the node feature matrix $X$ based on the connectivity defined by $A$. This function captures the relational information between nodes and can be defined as:
$f(X,A)\coloneqq\sigma_{\text{NN}}(D^{-\frac{1}{2}}(A+I)D^{-\frac{1}{2}}XW)$ (24)
where $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix, $I\in\mathbb{R}^{N\times N}$ is the identity matrix, $D\in\mathbb{R}^{N\times N}$ is the degree matrix of $A+I$, $X\in\mathbb{R}^{N\times D}$ is the node feature matrix, $W\in\mathbb{R}^{D\times W^{\prime}}$ is the layer’s weights where $W^{\prime}$ is the number of output features and $\sigma_{\text{NN}}(\cdot)$ is the activation function (e.g. ReLU). This results in an embedded vector for each node, $h_{i}\in\mathbb{R}^{W^{\prime}}$. In this work, three convolution layers are used where the embedded vector at each node, $h_{i}$, is the input of the next layer. This is described mathematically as:
$H_{1}$ $\coloneqq f_{W_{1}}(X,A)$ (25)
$H_{2}$ $\coloneqq f_{W_{2}}(H_{1},A)$ (26)
$H_{3}$ $\coloneqq f_{W_{3}}(H_{2},A)$ (27)
where $H_{1}\in\mathbb{R}^{N\times W_{1}^{\prime}},H_{2}\in\mathbb{R}^{N\times W_{2}^{\prime}},H_{3}\in\mathbb{R}^{N\times W_{3}^{\prime}}$ are the embedded node matrices at layers 1,2 and 3 and $W_{1},W_{2},W_{3}$ are the weight matrices that parameterize each layer. The terms $W_{1}^{\prime},W_{2}^{\prime},W_{3}^{\prime}\in\mathbb{Z}^{+}$ are defined as positive integers representing the number of output features for each layer respectively. These values are hyperparameters that determine the dimensionality of the embedded vectors produced by each layer. The Graph Convolutional Network (GCN), which is our graph module in the framework is used to learn hidden features that capture the structural information of the graph is shown in Figure 5.
Figure 5: Illustration of the Graph Convolutional Network (GCN) architecture, representing the Graph Module of our framework. The GCN takes the adjacency matrix and node feature matrix as input, processes them through three convolutional layers, each with ReLU activation, and outputs the embedded node matrix representing the learned representations of the nodes.
The resulting latent representations are then passed to our centralized value function $V_{\phi}(s)$, a fully connected network parameterized by $\phi$. This function assesses the value of actions, aiding in variance reduction of the policy $\pi_{\theta}$, another fully connected network parameterized by $\theta$, which maps states to probability density functions of actions. To train the policies, we employ Multi-Agent Proximal Policy Optimization (MAPPO). Mini-batches are sampled, and the objective is to minimize the loss function for each batch shown in Equation 2.3 which is used to update the parameters, $\theta$ and $\phi$.
A challenge of this method is that as the number of agents increases, the dimensionality of the node feature changes. To address this, we propose integrating global mean pooling within our framework, ensuring dimensionality remains constant as the number of agents increases.
Pooling operators in graphs were inspired by pooling methods in Convolutional Neural Networks (CNNs). Instead of simply concatenating all the latent hidden representations and feeding them directly into our central value function, we employ the global mean pool operator. The global mean pool operator aggregates information across all nodes in a graph by taking the average of the node features, providing a global representation of the graph without bias toward any specific node or agent. This approach ensures that the representation remains consistent, maintaining dimensionality across the fixed number of agents in the supply chain. The GNN outputs hidden feature matrix, $H\in\mathbb{R}^{N\times W_{3}^{\prime}}$ , can be denoted as $H=\{h_{1},h_{2},\dots,h_{N}\}$ where $N$ is the number of nodes and $h_{i}$ is the GNN output feature vector for node $i$. The global mean pooling operator computes the mean of each feature across all nodes.
Global Mean $\text{Global Mean Pool}(H)=\frac{1}{N}\sum_{i=1}^{N}h_{i}$ (28)
Alternatives to global mean pooling, including global max pooling, which could ignore contributions from “less active” nodes, potentially overlooking useful information from nodes that play a less prominent role in the system. Attention-based pooling dynamically weighs the importance of each node, which allows for a more flexible representation that can adapt to changing system dynamics. However, this introduces additional computational complexity and requires learning more parameters, increasing the risk of overfitting and making the model harder to optimize. The reader is referred to Grattarola et al. (2022) and Liu et al. (2022a) for further reading on different pooling strategies.
This effectively reduces the input dimensions to our critic from $N\times W_{3}^{\prime}$ to $1\times W_{3}^{\prime}$. While this dimensionality reduction improves computational efficiency, there is a potential drawback: it may result in the loss of specific spatial information. However, since each individual actor still leverages local information for each node, the loss is confined to the critic, whose role is mainly to provide global information. Moreover, the critic might not require the full level of spatial detail to assess the overall value of a state. The high-level features extracted through the global mean pooling mechanism might suffice. This idea has been explored in several studies such as Fujimoto et al. (2018) which highlights the critic’s focus on a good enough value approximation for policy improvement, suggesting it may not need full state representation, and Lyu et al. (2023) which proved that centralized critics may not be beneficial; particularly state-based critics can introduced unexpected bias. This aligns with the idea that the critic might not need a perfect representation but an accurate enough estimate for policy guidance. Finally, global mean pooling mechanisms can mitigate the risk of overfitting, a common problem in RL where the model performs well on training data but poorly on unseen data. By reducing the number of features, the model focuses on the most relevant aspects, which can lead to better generalization.
Figure 6 illustrates the framework which integrates Graph Convolutional Networks (GCNs) with Multi-Agent Proximal Policy Optimization (MAPPO) to enhance coordination among agents in a structured environment. The system is composed of three key components: the graph-based representation, the centralized critic, and the actors.
The environment is modeled as a graph, where nodes represent individual agents, and edges capture interactions between them as described in Section 3.2. Each agent’s local state information is encoded in a node feature matrix, while the adjacency matrix defines connectivity based on the edges (i.e., based on what supply chain nodes are connected). A Graph Convolutional Network (GCN) processes this structured input to extract informative representations by aggregating features from neighboring nodes.
The output of the GCN undergoes a global mean pooling operation, producing a single aggregated representation of the system’s state. This representation serves as input to the centralized critic, which estimates a shared value function to guide policy optimization. By leveraging this centralized value estimation, the approach facilitates effective coordination while maintaining decentralized execution.
Each agent is equipped with an actor network that receives local state information and outputs an action. These policies are trained using MAPPO, where the critic provides a learning signal to optimize the policies while maintaining stability in multi-agent settings as described in Section 2.3. During training, both the actor and critic networks are updated, with the critic leveraging the pooled global state to inform policy learning. However, during execution, only the decentralized actor networks are utilized, ensuring scalability and real-time decision-making without reliance on a central entity.
This architecture effectively combines graph-based feature extraction with multi-agent reinforcement learning, enabling agents to capture relational dependencies while optimizing their policies in a structured and scalable manner.
3.4 Regularized Graph Convolutional Networks (GCNs) combined with Multi-Agent Proximal Policy Optimization (MAPPO) and a Pooling strategy (Reg-P-GCN-MAPPO)
The problem of policy overfitting occurs as the estimated advantage is the same for all agents leading to the lack of credit assignment as described in Section 2.3. Theoretically, this can be solved by ensuring the shared advantage value does not affect other agents by decomposing the centralized advantage value for each agent. However, given the nature of multi-agent systems, this is impractical. Therefore, Gaussian noise can be introduced as a regularization technique to reduce bias in the advantage values. This approach has been widely adopted in various machine learning applications for its simplicity and effectiveness in regularization Goodfellow (2016); Igl et al. (2019).
Noise has been extensively studied and used in several forms across different components of reinforcement learning, including action space exploration and observation perturbation. In this work, we focus on introducing noise specifically in the value function which then propagates to the estimation of advantage values as shown in Figure 7. It is important to note that the framework shown in Figure 7 follows the same methodology described in Section 3.3, with the key distinction being the incorporation of noise in the value function. This approach offers several potential benefits and implications:
  1. 1.
  2. 2.
We sample a Gaussian noise,
$\epsilon_{n}\sim\mathcal{N}(0,\sigma^{2})$ (29)
$V_{\phi}(s)\leftarrow V_{\phi}(s)+\epsilon_{n}$ (30)
@@ 4 Results and Discussion
In this section four case studies of different inventory management configurations are used to illustrate the effectiveness of our proposed methodology as shown in Figure 8. Detailed descriptions of these supply chains can be found in C. We compare our proposed methodology (Reg-P-GCN-MAPPO and P-GCN-MAPPO) against other MARL methods including: IPPO, MAPPO, and Graph-based-MAPPO(G-MAPPO). This allows us to compare against state-of-the-art as well as analyze the contribution of the different components of our method. This is further benchmarked against single agent RL (PPO specifically) and a heuristic (s,S) policy. In the later, a static heuristic policy is found for each node in the network, where optimal parameters are found using a derivative-free method, implemented using SciPy (Virtanen et al., 2020) with a multi-start approach combined with local search.
All variant MARL algorithms are implemented in the Ray RLLib framework (Liang et al., 2018), and all the implementation details including neural network architectures and hyperparameters can be found in A and B.
@@ 4.1 Execution
Firstly, the bar graph shown in Figure 9 compares the cumulative profit for the different methodologies in the different supply chain network configurations. Notably, the proposed methodology, Reg-P-GCN-MAPPO, has on par performance or outperforms the other methodologies. This advantage becomes increasingly pronounced as the number of agents rises. This trend suggests that Reg-P-GCN-MAPPO may be particularly effective in large-scale, multi-agent inventory control scenarios. The decrease in cumulative profit from the 18 agent to 24 agent configuration can be attributed to differences in number of retailers present and the specific environment configuration chosen. Figure 9 also shows that while the static $(s_{t_{\text{inv,s}}}^{i},S_{t_{\text{ord,s}}}^{i})$ policy achieves similar cumulative profit performance to Reg-P-GCN-MAPPO, it exhibits a significantly higher standard deviation. This variability suggests that the static heuristic is less stable, potentially hindering its robustness in situations with unexpected changes to the environment.
This distinction is crucial in real-world applications, where supply chain networks and other multi-agent systems inevitably encounter unexpected events. RL’s ability to maintain good performance in uncertain environments is one of the key advantages compared to traditional methods. Therefore, the results highlight the strength of RL methods, particularly Reg-P-GCN-MAPPO, as its lower standard deviation suggests greater potential for robustness and adaptability in uncertain environments.
@@
Our methodology is also compared with other MARL methods. The notable difference between them is the information available to the critic. One key challenge in MARL is balancing access to global information (mitigating non-stationarity) with scalability limitations of naively combining all agent states. Our proposed methodologies, P-G-MAPPO and Reg-P-GCN-MAPPO, address this by leveraging graph-based approaches and the inherent structural properties of the problem. This allows agents to exploit both local and global information effectively.
@@
Figure 11 shows how IPPO consistently outperforms MAPPO, and this performance gap widens with a greater number of agents. This motivated the development of smarter methods of information aggregation, by leveraging graph-based approaches, where information can be harnessed whilst not suffering from performance shortcomings. This highlights the need for smarter information aggregation. G-MAPPO, P-GCN-MAPPO and Reg-P-GCN-MAPPO incorporate Graph Neural Networks (GNNs) to enable communication and information sharing while maintaining focus on local details.
However, from Figure 11, the G-MAPPO method does not always outperform MAPPO and IPPO methods. This can be attributed to the excessive capture of information, leading to overfitting of the policies. Specifically, when too much information is incorporated, the policies tend to overfit the training data. This is consistent with prior research, where Nayak et al. (2023) highlighted how excessive information aggregation in multi-agent systems without proper regularization can lead to overfitting, stressing the importance of intelligent information aggregation strategies to scale reinforcement learning methods effectively. Wang et al. (2022) also showed MAPPO may suffer from insufficient exploration due to its reliance on batch-sampled experiences from a replay buffer, which can lead to policy overfitting In contrast, in environments where exploration is critical, IPPO’s independent nature may facilitate more effective exploration strategies, allowing agents to discover diverse policies that are less prone to overfitting (Wang et al., 2021).
@@
In light of this, P-GCN-MAPPO outperforms G-MAPPO, MAPPO and IPPO in all 4 case studies. This occurs as the integration of a pooling mechanism, can help aggregate information from different parts of the graph, reducing the dimensionality of the data acting as a regularizer and potentially reducing overfitting. The message passing inherent in graph neural networks allows for agents to inherently communicate with each other whilst the pooling mechanism prevents overfitting of the policies. Despite losing local information in the critic, the global mean pooling mechanism allows for global information to be captured in the central critic that helps reduce the variance in the individual actors that are trained with local information.
@@
Our final proposed methodology hypothesizes that the addition of noise into the value function, propagates variability into the advantage function potentially reducing overfitting. The execution curves shown in Figure 11 show the impact of adding noise to the value function on our methodology’s performance. The results show that with a small number of agents (6 agents), adding noise into the value function does not outperform P-GCN-MAPPO compared to when the number of agents is larger (24 agents). When the number of agents is 6, the overall state space is smaller compared to a large agent system. In larger agent systems, the addition of noise can also be seen as regularization, where the policies become more robust to the complex environment.
@@
Moreover, it is useful to compare the change in entropy during the training phase for each of the different methods. In the context of MARL, entropy quantifies the uncertainty or randomness in the policy’s action distribution and is calculated based on the RL definition as the negative expectation of the log probabilities of the policy’s action distribution (Sutton, 2018). During the training process, as the agents learn, the entropy in the policy distribution decreases as the agents learn better policies. Figure 12 shows that the rate of entropy decrease differs between the different methods. The addition of noise to the value function introduces controlled variability which propagates into the advantage estimates. This helps maintain a higher entropy for a longer period during training which is desirable as high entropy promotes exploration. During MARL training, agents need to balance exploring new actions and exploiting known good actions. By exploring a wider range of actions (due to high entropy), agents are more likely to discover better strategies and avoid getting stuck in sub-optimal solutions. This can be especially important in complex environments where the best course of action depends on what other agents are doing. Figure 12 shows that Reg-P-GCN-MAPPO, maintains a higher entropy throughout training, promoting exploration and reducing policy overfitting leading to superior performance. Such behavior is particularly advantageous in environments characterized by stochasticity, where partial observability, randomness in rewards or transitions, and non-stationary dynamics are present (Su and Lu, 2022). In these settings, a policy with higher entropy is more likely to remain flexible, allowing agents to adapt to the evolving state of the environment. We hypothesize the persistence of higher entropy even after convergence is from the stochasticity introduced by entropy regularization, which leads the agents’ policies to stabilize around distributions that retain a degree of randomness. This is particularly crucial in cooperative scenarios, where agents must balance individual and collective goals. The added complexity of multiple interacting agents necessitates that policies remain responsive to the actions of others, which can manifest as sustained entropy within the policy distributions (Su and Lu, 2022).
In MARL environments, the non-stationary nature of the system, driven by the simultaneous learning and adaptation of multiple agents, further highlights the need for flexible policies. Each agent’s policy can influence the others, leading to a continuously changing environment. Policies that maintain higher entropy enable agents to remain adaptable, preventing them from becoming overly specialized or rigid in their strategies. This flexibility is crucial for sustaining performance in environments where the actions of other agents significantly impact the state dynamics (Nguyen et al., 2020; Li et al., 2020). Thus, while traditional methods may favor policies with lower entropy—indicative of more deterministic, greedy behaviors—the benefits of maintaining higher entropy are especially apparent in dynamic, multi-agent settings, where adaptability and flexibility are key to successful coordination and performance.
Figure 12: Entropy, as defined in reinforcement learning, represents the uncertainty or randomness in the policy’s action distribution (Sutton, 2018). The figure shows the change in entropy during the training process for the different methods in the 18 agent configuration